# High-dimensional sliced inverse regression with endogeneity

### v1

@@ 2.1 Review of sliced inverse regression and Lasso SIR
and $\mathbf{e}_{c}$ denotes the $c\times 1$ vector of ones. We can then estimate $\bm{\Lambda}^{*}$ by $\widehat{\bm{\Lambda}}^{*}=\widehat{\bm{\Sigma}}_{\mathbf{X}}-\widehat{\mathbf{T}}=n^{-1}\mathcal{X}^{\top}\mathbf{D}\mathcal{X}$, where $\mathbf{D}=\mathbf{I}_{n}-\{c/(c-1)\}(\mathbf{I}_{H}\otimes\mathbf{P}_{c})$. It follows that the eigenvector $\bm{\eta}^{*}_{k}$ can be estimated by $\hat{\bm{\eta}}^{*}_{k}$, the eigenvector associated with the $k$th largest eigenvalue $\hat{\lambda}^{*}_{k}$ of $\widehat{\bm{\Lambda}}^{*}$.
@@
For settings where $p$ is large and $\bm{\beta}_{k}$ is sparse, consider replacing the population relationship $\bm{\Sigma}_{\mathbf{X}}\bm{\beta}_{k}\propto\bm{\eta}^{*}_{k}$ by its sample counterpart. Then we have $\mathcal{X}^{\top}\mathcal{X}\bm{\beta}_{k}\propto\hat{\bm{\eta}}^{*}_{k}=(\hat{\lambda}^{*}_{k})^{-1}\mathcal{X}^{\top}\mathbf{D}\mathcal{X}\hat{\bm{\eta}}^{*}_{k}$, from which Lin et al., (2019) proposed the Lasso SIR estimator
$\hat{\bm{\beta}}_{k}^{*}=\arg\!\min_{\bm{\beta}}\frac{1}{2n}\left\lVert\tilde{\bm{y}}^{*}_{k}-\mathcal{X}\bm{\beta}\right\rVert_{2}^{2}+\mu_{k}\|\bm{\beta}\|_{1},$ (3)
@@
with the pseudo-response vector $\tilde{\bm{y}}^{*}_{k}=(\hat{\lambda}^{*}_{k})^{-1}\mathbf{D}\mathcal{X}\hat{\bm{\eta}}^{*}_{k}$, and $\mu_{k}>0$ denotes a tuning parameter for the $k$th pseudo-response. The Lasso penalty can be straightforwardly replaced by a more general sparsity-inducing penalty $\sum_{j=1}^{p}P_{n}(|\beta_{j}|)$ (e.g., the adaptive Lasso penalty in Nghiem et al., 2023c, ). With an appropriate choice of $\mu_{k}$, the solution $\hat{\bm{\beta}}_{k}$ is sparse, facilitating variable selection for each dimension of the estimator. Variable selection for the estimated central subspace is obtained by taking the union of the estimated support sets across all dimensions $k=1,\ldots,d$.
@@ 2.2 Inconsistency of SIR under endogeneity
When endogeneity is present, the inverse regression curve $E(\mathbf{X}\mid y)$ in no longer guaranteed to lie in the subspace spanned by $\bm{\Sigma}_{X}\mathbf{B}$, even when Condition 2.1 holds for the marginal $\mathbf{X}$. The following proposition establishes this result in the case $(\mathbf{X},\varepsilon)^{\top}$ follows a multivariate Gaussian distribution.
@@
Assume $(\mathbf{X}^{\top},\varepsilon)^{\top}$ follows a $(p+1)-$dimensional Gaussian distribution. Let $\bm{\Sigma}_{\mathbf{X},\varepsilon}=\operatorname{\text{Cov}}(\mathbf{X},\varepsilon)$. Then the inverse regression curve $E(\mathbf{X}\mid y)$ is contained in the linear subspace spanned by $\bm{\Sigma}_{\mathbf{X}}\mathbf{B}$ if and only if $\bm{\Sigma}_{\mathbf{X}}^{-1}\bm{\Sigma}_{\mathbf{X},\varepsilon}\in\text{col}(\mathbf{B})$, where $\text{col}(\mathbf{B})$ denotes the column space of $\mathbf{B}$.
In Proposition 1 the matrix $\bm{\Sigma}_{\mathbf{X}}^{-1}\bm{\Sigma}_{\mathbf{X},\varepsilon}$ is the coefficient vector of the best linear predictor of $\varepsilon$ from $\mathbf{X}$, so this proposition implies that even when endogeneity is present, the population foundation for SIR only holds when the random noise can be predicted from the linear combination $\mathbf{B}^{\top}\mathbf{X}$ itself. This condition is obviously restrictive, since it requires the same set of linear combinations to be used to predict both the response $y$ and the random noise $\varepsilon$.
When the condition in Proposition 1 is not satisfied, the Lasso SIR estimator in (3) is generally not a consistent estimator of the column space of $\mathbf{B}$. The following proposition makes this precise in a simple setting where the number of covariates $p$ is fixed and the penalization is asymptotically negligible.
@@
Let $\widehat{\mathbf{B}}^{*}$ denote the Lasso SIR estimator formulated in (3) with tuning parameters $\mu_{k}\to 0$ for all $k=1,\ldots,d$ as $n\to\infty$. Then, $\mathcal{P}(\widehat{\mathbf{B}}^{*})\to\mathcal{P}(\mathbf{B}^{*})$, where $\mathbf{B}^{*}=\bm{\Sigma}_{X}^{-1}\bm{\eta}^{*}$, with $\bm{\eta}^{*}$ the eigenvectors of $\bm{\Lambda}^{*}=\operatorname{\text{Cov}}\left\{E(\mathbf{X}\mid y)\right\}$. As a result, if the covariance vector $\bm{\Sigma}_{\mathbf{X},\varepsilon}\notin\bm{\Sigma}_{\mathbf{X}}\text{col}(\mathbf{B})$, then $\widehat{\mathbf{B}}^{*}$ is inconsistent for the column space of $\mathbf{B}$, and $\|P(\widehat{\mathbf{B}}^{*})-P(\mathbf{B})\|\neq o_{p}(1)$ as $n\to\infty$.
Proposition 1 implies that when the covariance vector $\bm{\Sigma}_{\mathbf{X},\varepsilon}\notin\bm{\Sigma}_{\mathbf{X}}\text{col}(\mathbf{B})$, endogeneity invalidates SIR as an estimator of the overall central subspace. Note this does not necessarily imply selection inconsistency, i.e. inconsistency in estimating the support set $S=\{1,\ldots,s\}$ corresponding to the non-zero elements of the diagonal of $\mathcal{P}(\mathbf{B})$. Selection consistency can still be achieved if the column space of $\mathbf{B}^{*}$ has the same support as that of $\mathbf{B}$; see the Supplementary Material S1 for an empirical demonstration. On the other hand, when the number of covariates $p$ grows then it is unlikely $\mathbf{B}^{*}$ will share the same support as $\mathbf{B}$. This motivates us to propose a new SIR method to handle endogeneity in high dimensional settings.
@@ 3 A two-stage Lasso SIR estimator
To overcome the general inconsistency of the SIR estimator when endogeneity is present, we propose a two-stage approach for SDR and variable selection inspired by the classical two-stage least-squares estimator (Hansen et al.,, 2008). Let $\widehat{\mathbf{X}}=E(\mathbf{X}\mid\mathbf{Z})=\bm{\Gamma}^{\top}\mathbf{Z}$, the conditional expectation of the covariates on the instruments. We first impose a new linearity condition to ensure the conditional expectation $E(\widehat{\mathbf{X}}\mid y)$ behaves similarly to $E(\mathbf{X}\mid y)$ when no endogeneity is present.
Condition 3.1.
@@
Despite being stronger than Condition 2.1, Condition 3.1 is satisfied when the linear instrument variable model holds and the $p+q$-dimensional random vector $(\mathbf{X}^{\top},\mathbf{Z}^{\top})^{\top}$ follows an elliptical distribution. In our simulation study in Section 5, we show the proposed method still exhibits reasonable performance even when this condition is not satisfied, for example, when $\mathbf{Z}$ is binary. We then have the following result.
@@
Assume Condition (3.1) holds. Then the quantity $E(\widehat{\mathbf{X}}\mid y)$ is contained in the linear subspace spanned by $\bm{\Sigma}_{\widehat{\mathbf{X}}}\mathbf{B}$.
Note the covariance matrix $\bm{\Sigma}_{\widehat{\mathbf{X}}}$ need not be of full-rank and invertible: this occurs if the number of endogenous covariates $p$ is greater than the number of instruments $q$. Nevertheless, Lemma 1 motivates us to propose a two-stage Lasso SIR (abbreviated to 2SLSIR) estimator for the column space of $\mathbf{B}$, assuming its entries are sparse, based on the observed vector $\mathbf{y}$, covariate matrix $\mathcal{X}$, and instrument matrix $\mathcal{Z}$. As we shall see, while theoretically we need to understand the behavior of its inverse, computationally $\bm{\Sigma}_{\widehat{\mathbf{X}}}$ does not have to be inverted to construct the 2SLSIR estimator.
The two stages of the 2SLSIR estimator are detailed as follows.
  • Stage 1: Identify and estimate the effects of the instruments $\mathbf{Z}$, and subsequently obtain the predicted values of the covariates $\mathbf{X}$. To accomplish this, we fit a multivariate instrumental variable model given by the second equation in (2), with the covariates $\mathbf{X}$ being the $p$-dimensional response and instruments $\mathbf{Z}$ being the $q$-dimensional predictor. We assume each endogenous covariate depends only on a few instrumental variables (Belloni et al.,, 2012; Lin et al.,, 2015), thus motivating a sparse first-stage estimator $\hat{\bm{\Gamma}}$ that can be computed in parallel by computing the estimated $j$th column as
    $\hat{\bm{\gamma}}_{j}=\arg\!\min_{\bm{\gamma}}\dfrac{1}{2n}\|\mathbf{x}_{j}-\mathcal{Z}\bm{\gamma}\|_{F}^{2}+\mu_{1j}\|\bm{\gamma}\|_{1},\;j=1,\ldots,p,$
    where $\mu_{1j}>0$ denotes the tuning parameter for the $j$th covariate in the first stage. After the estimator $\hat{\bm{\Gamma}}=[\hat{\bm{\gamma}}_{1},\ldots,\hat{\bm{\gamma}}_{p}]$ is formed, we construct the first-stage prediction matrix $\widehat{\mathcal{X}}=\mathcal{Z}\widehat{\bm{\Gamma}}$.
  • Stage 2: Compute a sparse estimator for the column space of $\mathbf{B}$ by applying the Lasso SIR estimator (3) with the response vector $\mathbf{y}$ and the first-stage prediction matrix $\hat{\mathcal{X}}$, Specifically, let $\widehat{\bm{\Lambda}}=(nc)^{-1}\widehat{\mathcal{X}}^{\top}\mathbf{D}\widehat{\mathcal{X}}=(nc)^{-1}\widehat{\bm{\Gamma}}^{\top}{\mathcal{Z}}^{\top}\mathbf{D}\mathcal{Z}\widehat{\bm{\Gamma}}$, and define the pseudo-response vector $\tilde{\bm{y}}_{k}=(\hat{\lambda}_{k})^{-1}\mathbf{D}\widehat{\mathcal{X}}\hat{\bm{\eta}}_{k}$, where $\hat{\bm{\eta}}_{k}$ denotes the eigenvector associated with the $k$th largest eigenvalue $\hat{\lambda}_{k}$ of $\hat{\bm{\Lambda}}$, and $\mathbf{D}$ is defined as in Section 2.1. Then the 2SLSIR estimator for the $k$th column of $\mathbf{B}$ is given by
    $\hat{\bm{\beta}}_{k}=\arg\!\min_{\bm{\beta}}\dfrac{1}{2n}\|\tilde{\bm{y}}_{k}-\widehat{\mathcal{X}}\bm{\beta}\|_{2}^{2}+\mu_{2k}\|\bm{\beta}\|_{1},\;k=1,\ldots,d,$ (4)
@@
We highlight that, computationally, both stages can be implemented straightforwardly and in parallel using standard coordinate descent algorithms. In this paper, we apply the R package glmnet (Friedman et al.,, 2010) to implement the 2SLSIR estimator, and select the tuning parameters $\mu_{j}$ and $\alpha_{k}$ via default standard ten-fold cross-validation as available in the software.
@@ 4 Theoretical analysis
We study the properties of the 2SLSIR estimator for the column space of $\mathbf{B}$ under the multiple index model (2), when the dimensionality of the covariates $p$ and instruments $q$ can grow exponentially with sample size $n$. Compared with the theoretical results for the linear outcome model in Lin et al., (2015) say, the major challenge here lies in the nonlinear nature of the conditional expectation $E(\widehat{\mathbf{X}}\mid y)$, and subsequently the behavior of the eigenvectors corresponding to the covariance matrix $\bm{\Lambda}=\operatorname{\text{Cov}}\left\{E(\widehat{\mathbf{X}}\mid y)\right\}$ which require more complex technical developments. We will use $C$, $C^{\prime}$, $C^{\prime\prime}$ to denote generic positive constants. Also, without loss of generality, we assume each column of the instrumental matrix $\mathcal{Z}$ is standardized such that $\mathbf{z}_{j}^{\top}\mathbf{z}_{j}=1$ for $j=1,\ldots,q$.
Following the notation in Wainwright, (Chapter 7, 2019), we say an $n\times m$ matrix $\mathbf{A}$ satisfies a restricted eigenvalue (RE) condition over a set $S$ with parameters $(\kappa,\nu)$ if $\|\mathbf{A}^{1/2}\bm{\delta}\|_{2}^{2}\geq\kappa\|\bm{\delta}\|_{2}^{2},$ for all $\bm{\delta}\in\mathcal{C}(S,\nu)=\left\{\delta\in\mathbb{R}^{m}:~{}\|\bm{\delta}_{S^{c}}\|_{1}\leq\nu\|\bm{\delta}_{S}\|_{1}\right\}$. Also, for a vector $\mathbf{v}$ let $\text{supp}(\mathbf{v})=\{j:v_{j}\neq 0\}$ be the support set of $\mathbf{v}$. Let $\sigma_{\max}=\max_{1\leq j\leq p}\sigma_{j}$, where we recall $\sigma^{2}_{j}$ are the diagonal elements of the covariance matrix $\bm{\Sigma}$ corresponding to $(\mathbf{U}_{i}^{\top},\varepsilon_{i})^{\top}\in\mathbb{R}^{p+1}$. We assume each $\sigma_{j}^{2}$ and hence $\sigma_{\max}$ to be a constant with respect to sample size.
The following technical conditions on the distribution of the instruments $\mathbf{Z}$ and the conditional expectation $E(\mathbf{X}\mid\mathbf{Z})=\bm{\Gamma}^{\top}\mathbf{Z}$ are required.
Condition 4.1.
The instrument random vector $\mathbf{Z}$ is sub-Gaussian with a $q\times q$ covariance matrix $\bm{\Sigma}_{\mathbf{Z}}$, such that for any $\mathbf{v}\in\mathbb{R}^{q}$ there exists a constant $C$ such that $\left\lVert\mathbf{Z}\right\rVert_{\psi_{2}}\leq C$, where $\left\lVert\mathbf{Z}\right\rVert_{\psi_{2}}=\sup_{\mathbf{v}\in\mathbb{R}^{q},\left\lVert\mathbf{v}\right\rVert=1}\left\lVert\mathbf{v}^{\top}\mathbf{Z}\right\rVert_{\psi_{2}},$ and $\left\lVert u\right\rVert_{\psi_{2}}=\sup_{m}m^{-1/2}\left(E|u|^{m}\right)^{1/m}$.
Condition 4.2.
There exists a constant $L>0$ such that $\|\bm{\Gamma}\|_{1}\leq L$.
Condition 4.3.
@@
The sub-Gaussian condition is often imposed in high-dimensional data analysis (e.g. Tan et al.,, 2018; Wainwright,, 2019; Nghiem et al., 2023b, ). Note a direct consequence of Condition 4.1 is that the instrumental matrix $\mathcal{Z}$ satisfies the RE condition with parameter $(\kappa,3)$ with high probability when $n>Cr\log q$ for some finite $\kappa$ (Banerjee et al.,, 2014); we will condition on this event in our development below. Furthermore, letting $\widehat{\bm{\Sigma}}_{\mathbf{Z}}=\mathcal{Z}^{\top}\mathcal{Z}/n$ denote the sample covariance of $\mathcal{Z}$ data, then
with probability at least $1-\exp(-C^{\prime}\log q)$ (see Ravikumar et al.,, 2011; Tan et al.,, 2018, for analogous results). Condition 4.2 is also imposed in the theoretical analysis of the two-stage penalized least square methods in Lin et al., (2015), and is necessary to bound the estimation error from the first stage. Finally, in the theoretical analysis below we allow the number of covariates $p$ and instruments $q$, and the sparsity levels $s$ and $r$ to depend on the sample size $n$, subject to the rate of growth of these parameters dictated by Condition 4.3. As a special case, if $r$ and $s$ are bounded then $p$ and $q$ can grow exponentially with $n$.

### v2

@@ 2.1 Review of SIR and lasso SIR
and ${e}_{c}$ denotes a $c\times 1$ vector of ones. We can then estimate $\bm{\Lambda}^{*}$ by $\widehat{\bm{\Lambda}}^{*}=\widehat{\bm{\Sigma}}_{\mathbf{X}}-\widehat{{T}}=n^{-1}\mathcal{X}^{\top}\mathbf{D}\mathcal{X}$, where $\mathbf{D}=\mathbf{I}_{n}-\{c/(c-1)\}(\mathbf{I}_{H}\otimes\mathbf{P}_{c})$. It follows that the eigenvector $\bm{\eta}^{*}_{k}$ can be estimated by $\hat{\bm{\eta}}^{*}_{k}=\bm{\eta}_{k}(\widehat{\bm{\Lambda}}^{*})$, that is, the eigenvector associated with the $k$th largest eigenvalue $\hat{\lambda}^{*}_{k}=\lambda_{k}(\widehat{\bm{\Lambda}}^{*})$ of $\widehat{\bm{\Lambda}}^{*}$ for $k=1,\ldots,d$.
@@
For settings where $p$ is large and $\bm{\beta}_{k}$ is sparse, consider replacing the population relationship $\bm{\Sigma}_{\mathbf{X}}\bm{\beta}_{k}\propto\bm{\eta}^{*}_{k}$ by its sample counterpart. Then we have $\mathcal{X}^{\top}\mathcal{X}\bm{\beta}_{k}\propto\hat{\bm{\eta}}^{*}_{k}=(\hat{\lambda}^{*}_{k})^{-1}\mathcal{X}^{\top}\mathbf{D}\mathcal{X}\hat{\bm{\eta}}^{*}_{k}$, from which Lin et al., (2019) proposed the lasso SIR estimator as
$\hat{\bm{\beta}}_{k}^{*}=\arg\!\min_{\bm{\beta}}\frac{1}{2n}\lVert\tilde{{\mathbf{y}}}^{*}_{k}-\mathcal{X}\bm{\beta}\rVert_{2}^{2}+\mu_{k}\|{\beta}\|_{1},$ (3)
@@
where $\tilde{{\mathbf{y}}}^{*}_{k}=(\hat{\lambda}^{*}_{k})^{-1}\mathbf{D}\mathcal{X}\hat{\bm{\eta}}^{*}_{k}$ is a pseudo-response vector and $\mu_{k}>0$ denotes a tuning parameter for the $k$th pseudo-response. Note the lasso penalty can be replaced by a more general sparsity-inducing penalty (e.g., the adaptive lasso penalty in Nghiem et al., 2023c, ), but such generalizations are outside the scope of this article. More importantly, with an appropriate choice of $\mu_{k}$ the solution $\hat{\bm{\beta}}_{k}$ is sparse, thus facilitating variable selection for each dimension of the estimator. Finally, variable selection for the estimated dimension reduction subspace is obtained by taking the union of the estimated support sets across all dimensions $k=1,\ldots,d$.
@@ 2.2 Inconsistency of SIR under endogeneity
When endogeneity is present, the inverse regression curve $E(\mathbf{X}\mid y)$ is no longer guaranteed to lie in the subspace spanned by $\bm{\Sigma}_{X}\mathbf{B}$, even when Condition 1 holds for the marginal $\mathbf{X}$. The following proposition establishes this result in the case $(\mathbf{X},\varepsilon)^{\top}$ follows a multivariate Gaussian distribution.
@@
Assume $(\mathbf{X}^{\top},\varepsilon)^{\top}$ follows a $(p+1)$-dimensional Gaussian distribution, and let $\bm{\Sigma}_{\mathbf{X},\varepsilon}=\operatorname{\text{cov}}(\mathbf{X},\varepsilon)$. Then the inverse regression curve $E(\mathbf{X}\mid y)$ is contained in the linear subspace spanned by $\bm{\Sigma}_{\mathbf{X}}\mathbf{B}$ if and only if $\bm{\Sigma}_{\mathbf{X}}^{-1}\bm{\Sigma}_{\mathbf{X},\varepsilon}\in\text{col}(\mathbf{B})$, where $\text{col}(\mathbf{B})$ denotes the column space of $\mathbf{B}$.
In Proposition 1 the matrix $\bm{\Sigma}_{\mathbf{X}}^{-1}\bm{\Sigma}_{\mathbf{X},\varepsilon}$ is the coefficient vector of the best linear predictor of $\varepsilon$ from $\mathbf{X}$. Hence this proposition implies that, when endogeneity is present, the population foundation for SIR can only hold when the random noise can be predicted from the linear combination $\mathbf{B}^{\top}\mathbf{X}$ itself. This condition is obviously restrictive, as it requires the same set of linear combinations to be used to predict both the response $y$ and the random noise $\varepsilon$.
Unsurprisingly, when the condition in Proposition 1 is not satisfied, then the lasso SIR estimator in (3) is also generally not a consistent estimator of the column space of $\mathbf{B}$. The following proposition makes this precise in the simple setting where the number of covariates $p$ is fixed, and the penalization is asymptotically negligible.
@@
Let $\widehat{\mathbf{B}}^{*}$ denote the lasso SIR estimator formulated in (3), with tuning parameters $\mu_{k}\to 0$ for all $k=1,\ldots,d$ as $n\to\infty$. Then, $\mathcal{P}(\widehat{\mathbf{B}}^{*})\to\mathcal{P}(\mathbf{B}^{*})$, where $\mathbf{B}^{*}=\bm{\Sigma}_{X}^{-1}\bm{\eta}^{*}$, where $\bm{\eta}^{*}$ is a $p\times d$ matrix of eigenvectors of $\Lambda^{*}=\operatorname{\text{cov}}\left\{E(\mathbf{X}\mid y)\right\}$. It follows that if the covariance vector $\Sigma_{\mathbf{X},\varepsilon}\notin\bm{\Sigma}_{\mathbf{X}}\text{col}(\mathbf{B})$, then $\|P(\widehat{\mathbf{B}}^{*})-P(\mathbf{B})\|\neq o_{p}(1)$ as $n\to\infty$ and $\widehat{\mathbf{B}}^{*}$ is inconsistent for the column space of $\mathbf{B}$.
Propositions 1 and 2 imply that when the covariance vector $\bm{\Sigma}_{\mathbf{X},\varepsilon}\notin\bm{\Sigma}_{\mathbf{X}}\text{col}(\mathbf{B})$, then endogeneity invalidates (lasso) SIR as an estimator of the overall dimension reduction subspace. Note this does not necessarily imply selection inconsistency, i.e. inconsistency in estimating the support set $S=\{1,\ldots,s\}$. Indeed, selection consistency can still be achieved if the column space of $\mathbf{B}^{*}$ has the same support as that of $\mathbf{B}$; see Appendix A for an empirical demonstration. On the other hand, when the number of covariates $p$ diverges with $n$, then it is unlikely $\mathbf{B}^{*}$ will share the same support as $\mathbf{B}$. This motivates us to propose a new SIR method to handle endogeneity in high-dimensional settings.
@@ 3 A two-stage lasso SIR estimator
To overcome the inconsistency of the SIR estimator when endogeneity is present, we propose a two-stage approach for SDR and variable selection. Let $\widehat{\mathbf{X}}=E(\mathbf{X}\mid\mathbf{Z})=\bm{\Gamma}^{\top}\mathbf{Z}$ denote the conditional expectation of the covariates on the instruments. We first impose a new linearity condition to ensure the conditional expectation $E(\widehat{\mathbf{X}}\mid y)$ behaves similarly to $E(\mathbf{X}\mid y)$ when no endogeneity is present.
Condition 2.
@@
Despite being stronger than Condition 1, Condition 2 is satisfied when the linear instrumental variable model holds and the $p+q$-dimensional random vector $(\mathbf{X}^{\top},\mathbf{Z}^{\top})^{\top}$ follows an elliptical distribution. In our simulation study in Section 5, we show the proposed method can still exhibit reasonable performance even when this condition is not satisfied, for example, when $\mathbf{Z}$ is binary. We then have the following result.
@@
Assume Condition (2) holds. Then the quantity $E(\widehat{\mathbf{X}}\mid y)$ is contained in the linear subspace spanned by $\bm{\Sigma}_{\widehat{\mathbf{X}}}\mathbf{B}$.
The covariance matrix $\bm{\Sigma}_{\widehat{\mathbf{X}}}$ in Lemma 1 need not be of full-rank and invertible e.g., we permit the number of endogenous covariates $p$ to exceed the number of instruments $q$. To be precise, while theoretically we need to understand the behavior of its inverse, computationally $\bm{\Sigma}_{\widehat{\mathbf{X}}}$ does not have to be inverted in what follows.
Lemma 1 motivates a two-stage lasso SIR estimator for the column space of $\mathbf{B}$ based on the observed vector $\mathbf{y}$, covariate matrix $\mathcal{X}$, and instrument matrix $\mathcal{Z}$. The two stages are detailed as follows: In stage 1, we identify and estimate the effects of the instruments $\bm{\Gamma}$, and subsequently obtain the predicted values of the covariates $\mathbf{X}$. To accomplish this, we fit a multivariate instrumental variable model given by the second equation in (2), with the covariates $\mathbf{X}$ as the $p$-dimensional response and instruments $\mathbf{Z}$ as the $q$-dimensional predictor. Moreover, because we assume each endogenous covariate depends only on a few instrumental variables, then a sparse first stage estimator $\hat{\bm{\Gamma}}$ is obtained by computing in parallel for $j=1,\ldots,p$
$\hat{\bm{\gamma}}_{j}=\arg\!\min_{\bm{\gamma}}\dfrac{1}{2n}\|\mathbf{x}_{j}-\mathcal{Z}\bm{\gamma}\|_{F}^{2}+\mu_{1j}\|\bm{\gamma}\|_{1},$
where $\mu_{1j}>0$ denotes the tuning parameter for the $j$th covariate in the first stage. Afterwards, the estimator $\hat{\bm{\Gamma}}=[\hat{\bm{\gamma}}_{1},\ldots,\hat{\bm{\gamma}}_{p}]$ is formed along with the fitted value matrix $\widehat{\mathcal{X}}=\mathcal{Z}\widehat{\bm{\Gamma}}$.
[…]
$\hat{\bm{\beta}}_{k}=\arg\!\min_{\bm{\beta}}\dfrac{1}{2n}\|\tilde{{\mathbf{y}}}_{k}-\widehat{\mathcal{X}}\bm{\beta}\|_{2}^{2}+\mu_{2k}\|\bm{\beta}\|_{1},\;k=1,\ldots,d,$ (4)
@@
Computationally, both stages can be implemented straightforwardly using coordinate descent algorithms e.g., we apply the R package glmnet (Friedman et al.,, 2010) in both stages.
@@ 4 Theoretical analysis
We study the properties of the two-stage lasso SIR estimator for the column space of $\mathbf{B}$}, under the multiple index model (2) and when the dimensionality of the covariates $p$ and instruments $q$ can grow exponentially with sample size $n$. Compared with the theoretical results for the linear outcome model in Lin et al., (2015), the major challenge here lies in the nonlinear nature of the conditional expectation $E(\widehat{\mathbf{X}}\mid y)$, and subsequently the behavior of the eigenvectors corresponding to the covariance matrix $\Lambda=\operatorname{\text{cov}}\{E(\widehat{\mathbf{X}}\mid y)\}$ which require more complex technical developments. In what follows, we will use $C$, $^{\prime}, C′′C^{\prime\prime}$ to denote generic positive constants.
Following Wainwright, (Chapter 7, 2019), we say an $n\times m$ matrix $\mathbf{A}$ satisfies a restricted eigenvalue condition over a set $S$ with parameters $(\kappa,\nu)$ if $\|\mathbf{A}^{1/2}\bm{\delta}\|_{2}^{2}\geq\kappa\|\bm{\delta}\|_{2}^{2},$ for all $\bm{\delta}\in\mathcal{C}(S,\nu)=\left\{\delta\in\mathbb{R}^{m}:~\|\bm{\delta}_{S^{c}}\|_{1}\leq\nu\|\bm{\delta}_{S}\|_{1}\right\}$. Also, for a vector $\mathbf{v}$ let $\text{supp}(\mathbf{v})=\{j:v_{j}\neq 0\}$ denote the support set of $\mathbf{v}$, and $\sigma_{\max}=\max_{1\leq j\leq p}\sigma_{j}$ where we recall $\sigma^{2}_{j}$ are the diagonal elements of the covariance matrix $\Sigma$ corresponding to $(\mathbf{U}_{i}^{\top},\varepsilon_{i})^{\top}\in\mathbb{R}^{p+1}$. We assume each $\sigma_{j}^{2}$ and hence $\sigma_{\max}$ to be a constant with respect to $n$.
The following technical conditions on the distribution of the $\mathbf{Z}$ and the conditional expectation $E(\mathbf{X}\mid\mathbf{Z})=\bm{\Gamma}^{\top}\mathbf{Z}$ are required.
Condition 3.
The instrument random vector $\mathbf{Z}$ is sub-Gaussian with a $q\times q$ covariance matrix $\bm{\Sigma}_{\mathbf{Z}}$, and for any $\mathbf{v}\in\mathbb{R}^{q}$ there exists a constant $C$ such that $\lVert\mathbf{Z}\rVert_{\psi_{2}}\leq C$, where $\lVert\mathbf{Z}\rVert_{\psi_{2}}=\sup_{\mathbf{v}\in\mathbb{R}^{q},\lVert\mathbf{v}\rVert=1}\lVert\mathbf{v}^{\top}\mathbf{Z}\rVert_{\psi_{2}},$ and $\lVert u\rVert_{\psi_{2}}=\sup_{m}m^{-1/2}\left(E|u|^{m}\right)^{1/m}$.
Condition 4.
There exist positive constants $L$ and $C$ such that $\|\bm{\Gamma}\|_{1}\leq L$, and the largest singular values $\|\bm{\Gamma}\|_{2}\leq C$.
Condition 5.
@@
The sub-Gaussian condition is common in high-dimensional data analysis (e.g. Tan et al.,, 2018; Wainwright,, 2019; Nghiem et al., 2023b, ), with a direct consequence being that the instrumental matrix $\mathcal{Z}$ satisfies the restricted eigenvalue condition with parameter $(\kappa,3)$ with high probability when $n>Cr\log q$, for some finite $\kappa$ (Banerjee et al.,, 2014); we will condition on this event in our development below. Furthermore, letting $\widehat{\bm{\Sigma}}_{\mathbf{Z}}=\mathcal{Z}^{\top}\mathcal{Z}/n$ denote the sample covariance of $\mathcal{Z}$ data, then
$\|\widehat{\bm{\Sigma}}_{\mathbf{Z}}-\bm{\Sigma}_{\mathbf{Z}}\|_{\max}=C\sqrt{\frac{\log q}{n}},$ (5)
with probability at least $1-\exp(-C^{\prime}\log q)$ (see also Ravikumar et al.,, 2011; Tan et al.,, 2018). Condition 4 is also imposed in the theoretical analysis of the two-stage penalized least square methods in Lin et al., (2015), and is necessary to bound the estimation error from the first stage. Finally, we allow the number of covariates $p$, instruments $q$, and the sparsity levels $s$ and $r$ to depend on the sample size $n$, subject to the rate of growth of these parameters dictated by Condition 5. For example, if $r$ and $s$ are bounded, then $p$ and $q$ can grow exponentially with $n$.